OpenStudy (australopithecus):

Determine if the series \[\sum_{n=1}^{\infty}n/e^{n} \] is convergent using the comparison or integral test. I used the integral test and got -1 + 2/e by computing the integral. Did I select the right test and answer this question correctly?

OpenStudy (anonymous):

converges for sure

OpenStudy (anonymous):

and integral test works nicely here since \[\int\frac{xdx}{e^x}=-(x+1)e^{-x}\] and so \[\int_1^{\infty}\frac{xdx}{e^x}\] is finite